It took John 12 hours riding his bike to make the round trip to his uncles, if he averaged 20 miles per hour out and 30 miles per hour back, how long did he travel each way
Answer by Alan3354(69443)

It took John 12 hours riding his bike to make the round trip to his uncles, if he averaged 20 miles per hour out and 30 miles per hour back, how long did he travel each way
=========================
Do it like this one:
-----------
If you drove to Austin at 70 miles per hour and returned at 50 miles per hour.what is the distance driven to Austin if total driving time was 7.2 hours
-----------
Avg speed for the round trip = 2*70*50/(70+50)
= 7000/120
= 175/3 mi/hr
------------
d = r*t = (175/3)*7.2
d = 420 miles RT
one-way = 210 miles